OpenStudy (anonymous):

I think i have the answer, i just need someone to confirm...Find the value of (fg)'(2)!!

OpenStudy (ghazi):

where is the function?

OpenStudy (anonymous):

f(2)=4, f ' (2)= -2, g(2) = 5, and g ' (2) = 3

OpenStudy (ghazi):

what you got?

OpenStudy (amistre64):

think product rule

OpenStudy (anonymous):

(fg)' = f * g' + g * f ' = 4* 3 + 5 * (-2)

OpenStudy (amistre64):

as ghazi says; what do you have for us to confirm with?

OpenStudy (amistre64):

slow connection today :)

OpenStudy (amistre64):

thats correct

OpenStudy (amistre64):

f=4, f'=-2 g=5, g'= 3 makes it easier to see (fg)' = f' g+f g' =-2(5) + 4(3)

OpenStudy (anonymous):

thanks!